OpenStudy (yrivers36):

Solve: 1/x+4-1/x=1 I found the lcd but lost afterwards.

OpenStudy (anonymous):

do you mean 1/(x+4) - (1/x) = 1 ?

OpenStudy (yrivers36):

yes

OpenStudy (anonymous):

ok for this type of equation, it's easiest to get rid of the fractions. so you want to multiply both sides by (x+4)(x) to get rid of both fractions on the left side.

OpenStudy (yrivers36):

what happens to the 1

OpenStudy (anonymous):

well, it gets multiplied by the (x+4)(x) just like the left side.

OpenStudy (yrivers36):

oh ok

OpenStudy (anonymous):

so can you move forward from there or are you still stuck?

OpenStudy (yrivers36):

I came up with x-x+4=x(x=4)

OpenStudy (anonymous):

be careful with your parentheses, what you have there doesn't look quite right, but almost there

OpenStudy (yrivers36):

Ok thank you I'm going to see if i can go further

OpenStudy (anonymous):

ok let me know if you need more help. remember when you are done you can always plug your answer back into the equation to see if it works.